# RL for Consistency Models: Faster Reward Guided Text-to-Image Generation

### v1

@@ Abstract
Reinforcement learning (RL) has improved guided image generation with diffusion models by directly optimizing rewards that capture image quality, aesthetics, and instruction following capabilities. However, the resulting generative policies inherit the same iterative sampling process of diffusion models that causes slow generation. To overcome this limitation, consistency models proposed learning a new class of generative models that directly map noise to data, resulting in a model that can generate an image in as few as one sampling iteration. In this work, to optimize text-to-image generative models for task specific rewards and enable fast training and inference, we propose a framework for fine-tuning consistency models via RL. Our framework, called Reinforcement Learning for Consistency Model (RLCM), frames the iterative inference process of a consistency model as an RL procedure. RLCM improves upon RL fine-tuned diffusion models on text-to-image generation capabilities and trades computation during inference time for sample quality. Experimentally, we show that RLCM can adapt text-to-image consistency models to objectives that are challenging to express with prompting, such as image compressibility, and those derived from human feedback, such as aesthetic quality. Comparing to RL finetuned diffusion models, RLCM trains significantly faster, improves the quality of the generation measured under the reward objectives, and speeds up the inference procedure by generating high quality images with as few as two inference steps. Our code is available at https://rlcm.owenoertell.com.
@@ 1 Introduction
Diffusion models also suffer from slow inference since they must take many steps to produce competitive results. This leads to slow inference time and even slower training time. Even further, as a result of the number of steps we must take, the resultant Markov decision process (MDP) possesses a long time horizon which can be hard for RL algorithms optimize. To resolve this, we look to consistency models. These models directly map noise to data and typically require only a few steps to produce good looking results. Although these models can be used for single step inference, to generate high quality samples, there exits a few step iterative inference process which we focus on. Framing consistency model inference, instead of diffusion model inference, as an MDP (as shown in Fig. 2) admits a much shorter horizon. This enables faster RL training and allows for generating high quality images with just few step inference.
More formally, we propose a framework Reinforcement Learning for Consistency Models (RLCM), a framework that models the inference procedure of a consistency model as a multi-step Markov Decision Process, allowing one to fine-tune consistency models toward a downstream task using just a reward function. Algorithmically, we instantiate RLCM using a policy gradient algorithm as this allows for optimizing general reward functions that may not be differentiable. In experiments, we compare to the current more general method, DDPO (Black et al., 2024) which uses policy gradient methods to optimize a diffusion model. In particular, we show that on an array of tasks (compressibility, incompressibility, prompt image alignment, and LAION aesthetic score) proposed by DDPO, RLCM outperforms DDPO in most tasks in training time, inference time, and sample complexity (i.e., total reward of the learned policy vesus number of reward model queries used in training) (Section 5).
@@ 2 Related Works
Consistency models are another family of generative models which directly map noise to data via the consistency function (Song et al., 2023). Such a function provides faster inference generation as one can predict the image from randomly generated noise in a single step. Consistency models also offer a more fine-tuned trade-off between inference time and generation quality as one can run the multistep inference process (Algorithm 2, in Appendix A) which is described in detail in Section 3.2. Prior works have also focused on training the consistency function in latent space (Luo et al., 2023) which allows for large, high quality text-to-image consistency model generations. Sometimes, such generations are not aligned with the downstream for which they will be used. The remainder of this work will focus on aligning consistency models to fit downstream preferences, given a reward function.
@@
Popularized by Black et al. (2024); Fan et al. (2023), training diffusion models with reinforcement learning requires treating the diffusion model inference sequence as an Markov decision process. Then, by treating the score function as the policy and updating it with a modified PPO algorithm (Schulman et al., 2017), one can learn a policy (which in this case is a diffusion model) that optimizes for a given downstream reward. Further work on RL fine-tuning has looked into entropy regularized control to avoid reward hacking and maintain high quality images (Uehara et al., 2024). Another line of work uses deterministic policy gradient methods to directly optimize the reward function when the reward function is differentiable (Prabhudesai et al., 2023). Note that when reward function is differentiable, we can instantiate a deterministic policy gradient method in RLCM as well. We focus on REINFORCE (Williams, 1992) style policy gradient methods for the purpose of optimizing a black-box, non-differentiable reward functions.
@@ 3.1 Reinforcement Learning
We model our sequential decision process as a finite horizon Markov Decision Process (MDP), $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\mu,H)$. In this tuple, we define our state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, initial state distribution $\mu$, and horizon $H$. At each timestep $t$, the agent observes a state $s_{t}\in\mathcal{S}$, takes an action according to the policy $a_{t}\sim\pi(a_{t}|s_{t})$ and transitions to the next state $s_{t+1}\sim P(s_{t+1}|s_{t},a_{t})$. After $H$ timesteps, the agent produces a trajectory as a sequence of states and actions $\tau=\left(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H},a_{H}\right)$. Our objective is to learn a policy $\pi$ that maximizes the expected cumulative reward over trajectories sampled from $\pi$.
$\mathcal{J}_{RL}(\pi)=\mathbb{E}_{\tau\sim p(\cdot|\pi)}\left[\sum_{t=0}^{H}R(s_{t},a_{t})\right]$
@@ 3.2 Diffusion and Consistency Models
$\text{d}x_{t}=\left[\bm{\mu}(\bm{x}_{t},t)-\frac{1}{2}\bm{\sigma}(t)^{2}\nabla\log p_{t}(\bm{x}_{t})\right]\text{d}t$
@@
After a model is trained, one can then trade inference time for generation quality with the multi-step inference process given in Appendix A, Algorithm 2. At a high level, the multistep consistency sampling algorithm first partitions the probability flow into $H+1$ points ($T=\tau_{0}>\tau_{1}>\tau_{2}\ldots>\tau_{H}=\epsilon$). Given a sample $x_{T}\sim p_{T}$, it then applies the consistency function $f_{\theta}$ at $(x_{T},T)$ yielding $\widehat{\bm{x}}_{0}$. To further improve the quality of $\widehat{\bm{x}}_{0}$, one can add noises back to $\widehat{\bm{x}}_{0}$ using the equation $\widehat{\bm{x}}_{\tau_{n}}\leftarrow\widehat{\bm{x}}_{0}+\sqrt{\tau_{n}^{2}-\tau_{H}^{2}}\bm{z}$, and then again apply the consistency function to $(\widehat{\bm{x}}_{\tau_{n}},\tau_{n})$, getting $\widehat{\bm{x}}_{0}$. One can repeat this process for a few more steps until the quality of the generation is satisfied. For the remainder of this work, we will be referring to sampling with the multi-step procedure. We also provide more details when we introduce RLCM later.
@@ 3.3 Reinforcement Learning for Diffusion Models
where clipping is used to ensure that when we optimize $p_{\theta}$, the new policy stay close to $p_{\theta_{old}}$, a trick popularized by the well known algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017).
In diffusion models, horizon $H$ is usually set as 50 or greater and time $T$ is set as $1000$. A small step size is chosen for the ODE solver to minimize error, ensuring the generation of high-quality images as demonstrated by Ho et al. (2020). Due to the long horizon and sparse rewards, training diffusion models using reinforcement learning can be challenging.
@@ 4 Reinforcement Learning for Consistency Models
Figure 2: Consistency Model As MDP: In this instance, $H=3$. Here we first start at a randomly sampled noised state $s_{0}\sim\left(\mathcal{N}(0,I),\delta_{\tau_{0}},p(\bm{c})\right)$. We then follow the policy by first plugging the state into the the consistency model and then noising the image back to $\tau_{1}$. This gives us $a_{0}$ which, based off of the transition dynamics becomes $s_{1}$. We then transition from $s_{1}$ by applying $\pi(\cdot)$, which applies the consistency function to $\widehat{x}_{0}$ and then noises up to $\tau_{2}$. We repeat this process until we reach timestep $H$. To calculate the end of trajectory reward, we apply the consistency function one more time to get a final approximation of $\widehat{x}_{0}$ and apply the given reward function to this image.
Modeling the MDP such that the policy $\pi(s)\mathrel{\leavevmode\hbox{\set@color\hskip 3.8889pt\leavevmode\hbox{\set@color$=$}\raisebox{4.66875pt}{\leavevmode\hbox{\set@color$\scriptstyle\Delta$}}\hskip 3.8889pt}}f_{\theta}(\bm{x}_{\tau_{t}},\tau_{t},\bm{c})+Z$ instead of defining $\pi(\cdot)$ to be the consistency function itself has a major benefit in the fact that this gives us a stochastic policy instead of a deterministic one. This allows us to use a form of clipped importance sampling like Black et al. (2024) instead of a deterministic algorithm (e.g. DPG (Silver et al., 2014)) which we found to be unstable and in general is not unbiased. Thus a policy is made up of two parts: the consistency function and noising with Gaussian noises. The consistency function takes the form of the red arrows in Fig. 2 whereas the noise is the green arrows. In other words, our policy is a Gaussian policy whose mean is modeled by the consistency function $f_{\theta}$, and covariance being $(\tau_{t}^{2}-\epsilon^{2})\mathbf{I}$ (here $\mathbf{I}$ is an identity matrix). Notice that in accordance with the sampling procedure in Algorithm 2, we only noise part of the trajectory. Note that the final step of the trajectory is slightly different. In particular, to calculate the final reward, we just transition with the applying the consistency function (red/yellow arrrow) and obtain the final reward there.
@@
4:     $\bm{X}_{0}\leftarrow P(f_{\theta},H,\bm{c})$triangleright where $\bm{X}_{0}$ is the batch of images, $\bm{x}_{0}$
@@
$\nabla_{\theta}\mathbb{E}_{\mathcal{D}}\sum_{t=1}^{T}\left[\min\left\{r(\bm{x}_{0},\bm{c})\cdot\frac{\pi_{\theta_{i+1}}(a_{t}|s_{t})}{\pi_{\theta_{i}}(a_{t}|s_{t})},r(\bm{x}_{0},\bm{c})\cdot\texttt{clip}\left(\frac{\pi_{\theta_{i+1}}(a_{t}|s_{t})}{\pi_{\theta_{i}}(a_{t}|s_{t})},1-\varepsilon,1+\varepsilon\right)\right\}\right]$
@@ 5 Experiments
Figure 3: Qualitative Generations: Representative generations from the pretrained models, DDPO, and RLCM. Across all tasks, we see that RLCM does not compromise the image quality of the base model while being able to transform naturalistic images to be stylized artwork that maximizes an aesthetic score, removes background content to maximize compression, and generate images of animals in fictional scenarios like riding a bike to maximize prompt-alignment.
@@
The aesthetic task is based off of the LAION Aesthetic predictor (Schuhmann, 2022) which was trained on 176,000 human labels of aesthetic quality of images. This aesthetic predictor is a MLP on top of CLIP embeddings (Radford et al., 2021). The images which produce the highest reward are typically artwork. This task has a smaller set of 45 animals as prompts.
@@
Figure 6: Inference Time: Plots showing the inference performance as a function of time taken to generate. For each task, we evaluated the final checkpoint obtained after training and measured the average score across 100 trajectories at a given time budget on 1 NVIDIA RTX A6000 GPU. We report the mean and std across three seeds for every run. Note that for RLCM, we are able to achieve high scoring trajectories with a smaller inference time budget than DDPO.
@@
Following the sample complexity evaluation proposed in Black et al. (2024), we first compare DDPO and RLCM by measuring how fast they can learn based on the number of reward model queries. As shown in Fig. 4, RLCM has better performance on three out of four of our tested tasks. Note that for the prompt-to-image alignment task, the initial consistency model finetuned by RLCM has lower performance than the initial diffusion model trained by DDPO. RLCM is able to close the performance gap between the consistency and diffusion model through RL finetuning [^4] It is possible that this performance difference on the compression and incompression tasks are due to the consistency models default image being larger. However, in the prompt image alignment and aesthetic tasks, we resized the images before reward calculation.. Fig. 3 demonstrates that similar to DDPO, RLCM is able to train its respective generative model to adapt to various styles just with a reward signal without any additional data curation or supervised finetuning.

### v2

@@ Abstract
Reinforcement learning (RL) has improved guided image generation with diffusion models by directly optimizing rewards that capture image quality, aesthetics, and instruction following capabilities. However, the resulting generative policies inherit the same iterative sampling process of diffusion models that causes slow generation. To overcome this limitation, consistency models proposed learning a new class of generative models that directly map noise to data, resulting in a model that can generate an image in as few as one sampling iteration. In this work, to optimize text-to-image generative models for task specific rewards and enable fast training and inference, we propose a framework for fine-tuning consistency models via RL. Our framework, called Reinforcement Learning for Consistency Model (RLCM), frames the iterative inference process of a consistency model as an RL procedure. Comparing to RL finetuned diffusion models, RLCM trains significantly faster, improves the quality of the generation measured under the reward objectives, and speeds up the inference procedure by generating high quality images with as few as two inference steps. Experimentally, we show that RLCM can adapt text-to-image consistency models to objectives that are challenging to express with prompting, such as image compressibility, and those derived from human feedback, such as aesthetic quality. Our code is available at https://rlcm.owenoertell.com.
@@ 1 Introduction
Diffusion models also suffer from slow inference since they must take many steps to produce competitive results. This leads to slow inference time and even slower training time. Even further, as a result of the number of steps we must take, the resultant Markov decision process (MDP) possesses a long time horizon which can be hard for RL algorithms optimize. To resolve this, we look to consistency models. These models directly map noise to data and typically require only a few steps to produce good looking results. Although these models can be used for single step inference, to generate high quality samples, there exits a few step iterative inference process which we focus on. Framing consistency model inference, instead of diffusion model inference, as an MDP admits a much shorter horizon. This enables faster RL training and allows for generating high quality images with just few step inference.
More formally, we propose a framework Reinforcement Learning for Consistency Models (RLCM), a framework that models the inference procedure of a consistency model as a multi-step Markov Decision Process, allowing one to fine-tune consistency models toward a downstream task using just a reward function. Algorithmically, we instantiate RLCM using a policy gradient algorithm as this allows for optimizing general reward functions that may not be differentiable. In experiments, we compare to the current more general method, DDPO (Black et al., 2024) which uses policy gradient methods to optimize a diffusion model. In particular, we show that on an array of tasks (compressibility, incompressibility, prompt image alignment, and LAION aesthetic score) proposed by DDPO, RLCM outperforms DDPO in tested compression, incompression, and aesthetic tasks in training time, inference time, and sample complexity (i.e., total reward of the learned policy versus number of reward model queries used in training) (Section 5).
@@ 2 Related Works
Consistency models (Song et al., 2023) are another family of generative models which directly map noise to data via the consistency function . Such a function provides faster inference generation as one can predict the image from randomly generated noise in a single step. Consistency models also offer a more fine-tuned trade-off between inference time and generation quality as one can run the multistep inference process (Algorithm 2, in Appendix A) which is described in detail in Section 3.2. Prior works have also focused on training the consistency function in latent space (Luo et al., 2023) which allows for large, high quality text-to-image consistency model generations. Sometimes, such generations are not aligned with the downstream for which they will be used. The remainder of this work will focus on aligning consistency models to fit downstream preferences, given a reward function.
@@
Popularized by Black et al. (2024); Fan et al. (2023), training diffusion models with reinforcement learning requires treating the diffusion model inference sequence as a Markov decision process. Then, by treating the score function as the policy and updating it with a modified PPO algorithm (Schulman et al., 2017), one can learn a policy (which in this case is a diffusion model) that optimizes for a given downstream reward. Further work on RL fine-tuning has looked into entropy regularized control to avoid reward hacking and maintain high quality images (Uehara et al., 2024). Another line of work uses deterministic policy gradient methods to directly optimize the reward function when the reward function is differentiable (Prabhudesai et al., 2023). Note that when reward function is differentiable, we can instantiate a deterministic policy gradient method in RLCM as well. We focus on REINFORCE (Williams, 1992) style policy gradient methods for the purpose of optimizing a black-box, non-differentiable reward functions.
@@ 3 Preliminaries
We provide some preliminary information on reinforcement learning, diffusion and consistency models, and discuss the application of reinforcement learning to diffusion models. Also note that we abuse notation and use $t$ to mean one of two things: the timestep along the diffusion trajectory or the timestep corresponding to the RL trajectory.
@@ 3.1 Reinforcement Learning
We model our sequential decision process as a finite horizon Markov Decision Process (MDP), $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\mu,H)$. In this tuple, we define our state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, initial state distribution $\mu$, and horizon $H$. At each timestep $t$, the agent observes a state $s_{t}\in\mathcal{S}$, takes an action according to the policy $a_{t}\sim\pi(a_{t}|s_{t})$ and transitions to the next state $s_{t+1}\sim P(s_{t+1}|s_{t},a_{t})$. After $H$ timesteps, the agent produces a trajectory as a sequence of states and actions $\tau=\left(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H},a_{H}\right)$. Our objective is to learn a policy $\pi$ that maximizes the expected cumulative reward over trajectories sampled from $\pi$,
$\mathcal{J}_{RL}(\pi)=\mathbb{E}_{\tau\sim p(\cdot|\pi)}\left[\sum_{t=0}^{H}R(s_{t},a_{t})\right].$
@@ 3.2 Diffusion and Consistency Models
$\text{d}x_{t}=\left[\bm{\mu}(\bm{x}_{t},t)-\frac{1}{2}\bm{\sigma}(t)^{2}\nabla\log p_{t}(\bm{x}_{t})\right]\text{d}t.$
@@
After a model is trained, one can then trade inference time for generation quality with the multi-step inference process given in Appendix A, Algorithm 2. At a high level, the multistep consistency sampling algorithm first partitions the probability flow into $H+1$ points ($T=\tau_{0}>\tau_{1}>\tau_{2}\ldots>\tau_{H}=\epsilon$). Given a sample $x_{T}\sim p_{T}$, it then applies the consistency function $f_{\theta}$ at $(x_{T},T)$ yielding $\widehat{\bm{x}}_{0}$. To further improve the quality of $\widehat{\bm{x}}_{0}$, one can add noise ($\bm{x}\sim\mathcal{N}(0,1)$) back to $\widehat{\bm{x}}_{0}$ using the equation $\widehat{\bm{x}}_{\tau_{n}}\leftarrow\widehat{\bm{x}}_{0}+\sqrt{\tau_{n}^{2}-\tau_{H}^{2}}\bm{z}$, and then again apply the consistency function to $(\widehat{\bm{x}}_{\tau_{n}},\tau_{n})$, getting $\widehat{\bm{x}}_{0}$. One can repeat this process for a few more steps until the quality of the generation is satisfied. For the remainder of this work, we will be referring to sampling with the multi-step procedure. We also provide more details when we introduce RLCM later.
@@ 3.3 Reinforcement Learning for Diffusion Models
where clipping is used to ensure that when we optimize $p_{\theta}$, the new policy stay close to $p_{\theta_{old}}$, a trick popularized by the well known algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). However, one could easily replace this with other policy gradient optimizers like Gao et al. (2024).
In diffusion models (and in our experiments for DDPO), horizon $H$ is usually set as 50 or greater and time $T$ is set as $1000$. A small step size is chosen for the ODE solver to minimize error, ensuring the generation of high-quality images as demonstrated by Ho et al. (2020). Due to the long horizon and sparse rewards, training diffusion models using reinforcement learning can be challenging.
@@ 4 Reinforcement Learning for Consistency Models
Figure 2: Consistency Model As MDP: In this instance, $H=3$. Here we first start at a randomly sampled noised state $s_{0}\sim\left(\mathcal{N}(0,I),\delta_{\tau_{0}},p(\bm{c})\right)$. We then follow the policy by first plugging the state into the the consistency model (red line) and then noising the image back to $\tau_{1}$ (green line). This gives us $a_{0}$ which, based off of the transition dynamics becomes $s_{1}$ (green circle). We then transition from $s_{1}$ by applying $\pi(\cdot)$, which applies the consistency function to $\widehat{x}_{0}$ and then noises up to $\tau_{2}$. To calculate the end of trajectory reward, we apply the consistency function one more time (yellow line) to get a final approximation of $\widehat{x}_{0}$ and apply the given reward function to this image. Note that the red and green lines on both sides of the diagram represent the same thing.
Modeling the MDP such that the policy $\pi(s)\mathrel{\leavevmode\hbox{\set@color\hskip 3.8889pt\leavevmode\hbox{\set@color$=$}\raisebox{4.66875pt}{\leavevmode\hbox{\set@color$\scriptstyle\Delta$}}\hskip 3.8889pt}}f_{\theta}(\bm{x}_{\tau_{t}},\tau_{t},\bm{c})+Z$ instead of defining $\pi(\cdot)$ to be the consistency function itself has a major benefit in the fact that this gives us a stochastic policy instead of a deterministic one. This allows us to use a form of clipped importance sampling like Black et al. (2024) instead of a deterministic algorithm (e.g. DPG (Silver et al., 2014)) which we found to be unstable and in general is not unbiased. Thus a policy is made up of two parts: the consistency function and noising with Gaussian noises. The consistency function takes the form of the red arrows in Fig. 2 whereas the noise is the green arrows. In other words, our policy is a Gaussian policy whose mean is modeled by the consistency function $f_{\theta}$, and covariance being $(\tau_{t}^{2}-\epsilon^{2})\mathbf{I}$ (here $\mathbf{I}$ is an identity matrix). Notice that in accordance with the sampling procedure in Algorithm 2, we only noise part of the trajectory. Note that the final step of the trajectory is slightly different. In particular, to calculate the final reward, we just apply the consistency function (red/yellow arrrow) and obtain the final reward.
@@
4:     $\bm{x}_{0}\leftarrow P(f_{\theta},H,\bm{c})$triangleright where $\bm{x}_{0}$ is the batch of images
@@
$\nabla_{\theta}\left[\min\left\{r({\bm{x}_{0,m}},\bm{c})\cdot\frac{\pi_{\theta_{m+1}}(a_{t}|s_{t})}{\pi_{\theta_{m}}(a_{t}|s_{t})},r({\bm{x}_{0,m}},\bm{c})\cdot\texttt{clip}\left(\frac{\pi_{\theta_{m+1}}(a_{t}|s_{t})}{\pi_{\theta_{m}}(a_{t}|s_{t})},1-\varepsilon,1+\varepsilon\right)\right\}\right]$
@@ 5 Experiments
Figure 3: Qualitative Generations: Representative generations from the pretrained models, DDPO, and RLCM. Across all tasks, we see that RLCM is able to finetune output of the model to fit specific reward functions. Due to the lack of regularization to the pretrained model, some artifacts (seen in the compression task) and significant similarity in output are indeed seen).
@@
The aesthetic task is based off of the LAION Aesthetic predictor (Schumman, 2022) which was trained on 176,000 human labels of aesthetic quality of images. This aesthetic predictor is a MLP on top of CLIP embeddings (Radford et al., 2021). The images which produce the highest reward are typically artwork. This task has a smaller set of 45 animals as prompts.
@@
Figure 6: Inference Time: Plots showing the inference performance as a function of time taken to generate. For each task, we evaluated the final checkpoint obtained after training and measured the average score across 100 trajectories at a given time budget on 1 NVIDIA RTX A6000 GPU. We report the mean and std across three seeds for every run. Note that for RLCM, we are able to achieve high scoring trajectories with a smaller inference time budget than DDPO. Final reward values may differ from previous plots due to random selection of prompts used for measurement.
@@
Following the sample complexity evaluation proposed in Black et al. (2024), we first compare DDPO and RLCM by measuring how fast they can learn based on the number of reward model queries. As shown in Fig. 4, RLCM has better performance on three out of four of our tested tasks in terms of number of reward queries. Note that for the prompt-to-image alignment task, the initial consistency model finetuned by RLCM has lower performance than the initial diffusion model trained by DDPO. RLCM is able to close the performance gap between the consistency and diffusion model through RL finetuning [^4] It is possible that this performance difference on the compression and incompression tasks are due to the consistency models default image being larger. However, in the prompt image alignment and aesthetic tasks, we resized the images before reward calculation.. Fig. 3 demonstrates that similar to DDPO, RLCM is able to train its respective generative model to adapt to various styles just with a reward signal without any additional data curation or supervised finetuning.
@@ 5.4 Qualitative Effects on Generalization
5.5 Convergence Results of Tasks
To compare fairly to Black et al. (2024), we only train for only the same number of reward queries which means that in two tasks (Aesthetic and Prompt Image Alignment) convergence of the tasks is not shown.
We trained DDPO and RLCM for longer on the aesthetic task and observed that RLCM asymptotically arrived at the approximate maximum reward value (value 10 is the maximum reward available in the training dataset for the reward model). For DDPO, when it runs longer (after 72 hours), it reaches a reward around 9.5, but unfortunately crashes.
We also attempted to run the text-image alignment task longer for DDPO, unfortunately we observed the same crashing behavior. We suspect that it is due to the fixed learning rate schedule used in the original DDPO codebase (note that for fair comparison, we use the original DDPO codebase with the default hyperparameters proposed by the authors of DDPO). Applying strategies like learning rate decay may stabilize DDPO, but it would require additional hyperparameter tuning for DDPO.
5.6 Known Limitations
The main known limitation observed throughout the use of RLCM is overfitting to the reward function. Indeed, as seen in Fig. 3, unrealistic generations as seen in the compression task or extremely similar backgrounds like in the aesthetic task do arise. In cases where such overfitting is undesirable, a KL regularized loss which incorporates some measure of divergence between the currently trained model and the initial model will improve generations. However, this was not a focus of this work.